# Immiscible Diffusion: Accelerating Diffusion Training with Noise Assignment

### v1

@@ 4.1 Experiment Settings
To elaborate the performance of Immiscible Diffusion, we utilize the proposed method on Consistency Models [38], DDIM [36] and Stable Diffusion [32], and using CIFAR-10 [14], CelebA [21], tiny and random picked 10% ImageNet [3] datasets due to the limitation of computation resource. The training hyperparameters are shown in Tab. 1. Unspecified hyperparameters are taken the same as those in their baseline methods’ original papers. For evaluations, we compare the results generated by our Immiscible Diffusion method and the baseline using both qualitative assessments and the quantitative evaluation metric FID [7].
Note that for Consistency Models, we use the single step generation consistency training. For DDIM, we add no noise during sampling and use linear scheduling for picking sampling steps. For Stable Diffusion, we directly use the implementation from Diffusers of Huggingface team [31].
@@
Model Consistency Model DDIM Consistency Model Stable Diffusion Unconditional Consistency Model Stable Diffusion Class-conditional
Dataset CIFAR-10 CIFAR-10 CelebA 10% ImageNet Tiny ImageNet ImageNet
Batch Size 512 256 1024 512 2048 2048
Resolution $32\times 32$ $32\times 32$ $64\times 64$ $256\times 256$ $64\times 64$ $256\times 256$
Devices $60004\times A$ 6000 $50001\times A$ 5000 $8008\times A$ 800 $60004\times A$ 6000 $80016\times A$ 800 $8008\times A$ 800
4.2 Main Results on Training Efficiency
@@
In Fig. 4, we show the FIDs of images generated with baseline and immiscible consistency models trained with different training steps on the CIFAR-10 dataset, the CelebA dataset and the tiny ImageNet dataset, respectively. We observe that the immiscible consistency model trains much faster than the baseline consistency model, and converges to a significantly lower FID on all of the CIFAR-10, CelebA, and tiny-ImageNet datasets. We also show the images generated by immiscible and baseline consistency models trained for 100k steps in Fig. 5, where we can see that the images generated by immiscible consistency model are much more complete and realistic. Tab. 2 further presents the training steps necessary to achieve specific reasonable FID thresholds. We find that the immiscible consistency model significantly improves the training efficiency by around 3x, proving the effectiveness of Immiscible Diffusion in training accelerations.
@@
In the main experiment, we observe that our method on top of the Consistency Model is effective across the datasets varying from different data size and resolutions. Indeed, the Consistency Model is a few-step diffusion model, and our proposed Immiscible Diffusion especially works on improving the denoising effect when the noise level is high, as shown in Fig. 3. The improvement of the training efficiency on such a few step diffusion model further validates our findings.
@@
One characteristic of the Consistency Model is that it approximates the SDE-diffusion model with ODE approximation. Thus the original image-noise mapping is highly jumbled together since it is highly possible that closed image data points are diffused to distant noise points. Our Immiscible Diffusion improves this issue by adjusting the trajectories of image-noise mapping and making them more distinguishable.
@@
Figure 4: Evaluation of baseline and immiscible consistency models on CIFAR-10, CelebA, and tiny-ImageNet dataset. We illustrate the FID of two models with different training steps. Clearly, immiscible consistency models have much higher efficiency than the traditional one.
Figure 5: Qualitative comparison for immiscible and baseline consistency model. We show images generated with the two models trained for 100k steps respectively. Compared to baseline method, immiscible models can catch more details and more feature of objects.
More baselines: To show the generalization of Immiscible Diffusion for more baselines, we further conduct experiments on two baselines: DDIM [36] and Stable Diffusion [30] on the CIFAR-10 and the randomly picked 10% ImageNet dataset, respectively. As shown in Fig. 6, we find that our immiscible DDIM significantly improves the training speed and the FID compared to those of the baseline DDIM on the CIFAR-10 dataset. This improvement demonstrates the effectiveness of our proposed method works beyond the Consistency Model and can be generalized to more few-step denoising models.
@@
Figure 6: Evaluation of baseline and immiscible DDIM on CIFAR-10 dataset, using 20 inference steps. (a) FID of two models with different training steps (b) Qualitative comparison of two models trained 100k steps.
@@
To further evaluate the generalizability on the popular baseline, Stable Diffusion [30], we conduct experiments on the ImageNet dataset. We observe that immiscible Stable Diffusion and baseline Stable Diffusion achieve similar FID without significant gap, yet our immiscible Stable Diffusion is able to generate much higher quality images from a subjective human judgement. For example, Fig. 7 shows our proposed method generates significantly clearer images compared to the baseline. More visualization without any cherry-picking can be seen in the supplementary materials. We indicate that even though FID is the primary metric and remarkably successful, the metric is known to sometimes disagree with human judgement [17].
@@
Class-conditional Generation: We extend immiscible diffusions to class-conditional generations on ImageNet dataset with Stable Diffusion [32], to explore the performance of immiscible diffusion in conditional cases. Results are shown in Fig. 8, where we observe that the FID for immiscible class-conditional Stable Diffusion is 20.90, which is 1.53 lower than our Stable Diffusion baseline. Qualitative comparisons further prove such performance enhancements, which augment the effectiveness of immiscible diffusions into more commonly-used conditional generations.
Figure 7: Images generated by immiscible and baseline Stable Diffusion trained on ImageNet for 70k steps. We see that the immiscible Stable Diffusion presents more reasonable modal and catch more general features and details.
@@ 4.3 Discussion
Figure 8: Evaluation of baseline and immiscible class-conditional Stable Diffusion on ImageNet dataset, using 20 inference steps. (a) FID of two models with different training steps (b) Qualitative comparison of two models trained 20k steps.
How much does image-noise distance reduce in the assignment? Tab. 3 shows the reduction in distance after the image-noise assignment. We find that the distance only reduces by about 2%, with a slight increase observed at higher batch sizes. However, this small change in distance is sufficient to effectively activate the denoising at high-noise levels, significantly boosting training efficiency. We attribute the low distance reduction rate after the assignment to the extremely high dimensionality (3072 for each image of the CIFAR-10 dataset) of the image and noise space. Furthermore, we indicate that our assignment method does not introduce significant extra overhead due to our utilization of quantized assignment in our practical implementation. Even for a large per-GPU batch size of 1024, our algorithm only brings in additional 22.8ms, demonstrating the potential of utilization for future applications.
Table 3: Image-noise datapoint L2 distance reduction after the assignment for minimizing it.
@@
Which distance function should be used for assignment? We use the L2 norm for our experiments. However, we note that the L2 norm may face more challenges in distance evaluation in high-dimensional spaces compared to the L1 norm. Therefore, we compare the performance of immiscible DDIMs using assignments based on the L1 and L2 norms. The results, as illustrated in Tab. 4, show that using the L2 norm provides better performance than the L1 norm.
Table 4: FID of using L1 or L2 norm for noise assginment in immiscible DDIM on CIFAR-10.
Traning Steps (k) 70.2 93.6 117.0 140.4 163.8
DDIM 6.30 5.56 4.86 4.34 4.12
Immiscible DDIM using L2 Norm 5.28 4.56 4.13 3.81 3.70
Immiscible DDIM using L1 Norm 5.34 4.66 4.16 3.87 3.82
How does Immiscible Diffusion work for different inference steps in DDIM? As we improve the utilization of high (noisy) diffusion steps, our performance improves, particularly when there are not too many redundant denoising layers. Specifically, we observe more significant performance boosts with immiscible DDIM when the number of inference denoising steps is lower, as shown in Tab. 5.
@@
$\Delta$FID -0.15 -0.17 -0.98 -1.49 -2.57
5 Conclusion, Limitations, and Future Work
Inspired by the immiscible phenomenon in physics, we introduce Immiscible Diffusion, a method to improve noise-data mapping to accelerate diffusion training. Specifically, Immiscible Diffusion is an assignment-then-diffusion strategy. It minimizes the image-noise pair distance within a mini-batch so that the similar images are diffused to nearby noise points. This simple approach requires only one line of code and includes a quantized-assignment strategy to reduce computational overhead. Experiments show our Immiscible Diffusion approach speeds up training by approximately 3x on the CIFAR-10 dataset, 1.3x on the CelebA dataset, and 1.2x on the tiny-ImageNet dataset. Thus, we show that Immiscible Diffusion can generalize to across datasets and different baselines. Further analysis is provided to explain how this works.
Limitation. Our current assignment strategy is straightforward and not necessarily optimal. Due to the limited computational resources, our experiments are mainly conducted on small-scale datasets, so we lack the validation on larger-scale dataset such as LAION. In future work, we will improve the assignment strategy to cater to practical utilization of conditional generation such as accelerating the text-to-image or text-to-video diffusion training.
Broader impact. With the increased use of diffusion models for image and video generation, the training of diffusion models is certain to become an increasing portion of data center workloads. Moreover, training time is significant bottleneck to model development. Our proposed method significantly improves the efficiency of diffusion model training. We believe our method has the potential to accelerate progress and to reduce the cost of development in this field.
Appendix A Supplemental material
A.1 Generated images from immiscible and baseline stable diffusion models trained on 10% ImageNet Dataset for 70k steps without cherry-picking
Figure 9: Generated images from immiscible and baseline stable diffusion models trained on 10% ImageNet Dataset for 70k steps without cherry-picking
A.2 Generated images from immiscible and baseline consistency models trained on CIFAR-10 Dataset for 100k steps without cherry-picking
@@ Appendix A Supplemental material
A.3 Generated images from immiscible and baseline consistency models trained on CelebA Dataset for 100k steps without cherry-picking

### v2

@@ 4.1 Experiment Settings
To elaborate the performance of Immiscible Diffusion, we utilize the proposed method on Consistency Models [47], DDIM [45] and Stable Diffusion [41], and using CIFAR-10 [20], CelebA [29], tiny, random picked 10% and the full ImageNet [4] datasets due to the limitation of computation resource. The training hyperparameters are shown in Tab. 1. Unspecified hyperparameters are taken the same as those in their baseline methods’ original papers. For evaluations, we compare the results generated by our Immiscible Diffusion method and the baseline using both the quantitative evaluation metric FID [10] and qualitative assessments.
Note that for Consistency Models, we use the single-step generation consistency training. For DDIM, we add no noise during the sampling and use linear scheduling for picking sampling steps. For Stable Diffusion, we directly use the implementation from Diffusers of Huggingface team [40]. For fine-tuning, we use Stable Diffusion v1.4 [40] as the pre-trained model.
@@
  Model Consistency Model Consistency Model Consistency Model DDIM Stable Diffusion Unconditional Stable Diffusion Class-conditional Stable Diffusion Fine-tuning
Dataset CIFAR-10 CelebA Tiny ImageNet CIFAR-10 10% ImageNet Full ImageNet Full ImageNet
Batch Size 512 1024 2048 256 512 2048 512
Resolution $32\times 32$ $64\times 64$ $64\times 64$ $32\times 32$ $256\times 256$ $256\times 256$ $256\times 256$
Devices $60004\times A$ 6000 $8008\times A$ 800 $80016\times A$ 800 $50001\times A$ 5000 $60004\times A$ 6000 $8008\times A$ 800 $60004\times A$ 6000
4.2 Training Efficiency Improvement with Linear Assignment
@@
Unconditional generation with Consistency Models: In Fig. 4, we show the FIDs of images generated with baseline and immiscible Consistency Models trained with different training steps on the CIFAR-10 dataset, the CelebA dataset and the tiny ImageNet dataset, respectively. We observe that the immiscible Consistency Model trains much faster than the baseline Consistency Model, and converges to a significantly lower FID on all these datasets. We also show the images generated by immiscible and baseline Consistency Models trained for 100k steps in Fig. 9 in the Supplemental Materials, where we find that the images generated by the Immiscible Consistency Model are much more complete and realistic. Tab. 3 in the Supplemental Materials further presents the training steps necessary to achieve specific reasonable FID thresholds. We find that the immiscible Consistency Model significantly improves the training efficiency by around 3x, proving the effectiveness of Immiscible Diffusion in training accelerations.
@@
In the main experiment, we observe that our method on top of the Consistency Model is effective across the datasets varying from different data sizes and resolutions. Indeed, the Consistency Model is a few-step diffusion model, and our proposed Immiscible Diffusion especially works on improving the denoising effect when the noise level is high, as shown in Fig. 3. The improvement of the training efficiency on such a few-step diffusion model further validates our findings.
@@
One characteristic of the Consistency Model is that it approximates the SDE-diffusion model with the ODE approximation. Thus, the original image-noise mapping is highly jumbled together since it is highly possible that closed image data points are diffused to distant noise points. Our Immiscible Diffusion improves this issue by adjusting the trajectories of image-noise mapping and making them more distinguishable.
@@
Figure 4: Evaluation of baseline and immiscible Consistency Models on (a) CIFAR-10, (b) CelebA, and (c) tiny-ImageNet dataset. We illustrate the FID of two models with different training steps. Clearly, immiscible Consistency Models have much higher efficiency than the vanilla ones.
@@
Figure 5: Evaluation of baseline and Immiscible DDIM on CIFAR-10 dataset with different inference steps $S$. We find that Immiscible DDIM ourperform the baseline more significantly when the number of inference steps $S$ is smaller.
@@
More baselines: To show the generalization of Immiscible Diffusion for more baselines, we further conduct experiments on two baselines: DDIM [45] and Stable Diffusion [39] on the CIFAR-10 and the randomly picked 10% ImageNet dataset, respectively. As shown in Fig. 5, and detailed in Tab. 4 in Supplemental Materials, we find that our immiscible DDIM significantly improves the training speed and the FID compared to those of the baseline DDIM on the CIFAR-10 dataset, and the improvement is more significant when the sampling step is lower. This demonstrates the effectiveness of our proposed method works beyond the Consistency Model and can be generalized to more few-step denoising models. We also provide a discussion in Part A.1.4, showing that the effectiveness of Immiscible Diffusion can persist in a wide range of batch sizes. To further evaluate generalizability on the popular baseline, Stable Diffusion [39], we also conduct unconditional generation experiments on the ImageNet dataset. We observe that immiscible Stable Diffusion and baseline Stable Diffusion achieve similar FID without significant gap, yet our immiscible Stable Diffusion is able to generate much higher quality images from a subjective human judgement. For example, Fig. 14 in the Supplemental Materials shows that our proposed method generates significantly clearer images compared to the baseline. More visualization without any cherry-picking can be seen in Fig. 15 in the supplementary materials. We indicate that even though FID is the primary metric and is remarkably successful, the metric is known to sometimes disagree with human judgement [23].
@@
Class-conditional Generation: We extend Immiscible Diffusion to class-conditional generations on ImageNet dataset with Stable Diffusion [41], to explore the performance of Immiscible Diffusion in conditional generations. Results are shown in Fig. 6 (a), where we observe that in 20k training steps, the FID for immiscible class-conditional Stable Diffusion is 16.43, which is 1.49 lower than our Stable Diffusion baseline. We further confirm such improvements on CMMD [15], where the immiscible and vanilla models get 1.385 and 1.436 respectively. Additional evaluation on CLIPScore [9] shows that both the immiscible and the baseline models generate images with CLIPScores of 28.55, with a standard deviation of 0.01 and 0.02 respectively, indicating that Immiscible Diffusion does not hurt the image-prompt correspondence in complicated ImageNet dataset. Qualitative comparisons in Fig. 16 in Supplemental Materials further prove such performance enhancements, which augment the effectiveness of Immiscible Diffusions into more commonly-used conditional generations.
Fine-tuning: Our Immiscible Diffusion can also be applied to enhance the fine-tuning process where numerous applications fall in. We fine-tune the stable diffusion v1.4 model [41] on ImageNet dataset, finding that immiscible fine-tuning achieves an FID of 10.28 compared to 11.45 for vanilla fine-tuning with 5k training steps. Detailed results are shown in Fig. 6. This further broadens the application of Immiscible Diffusion.
Figure 6: Evaluation of baseline and immiscible class-conditional Stable Diffusion on ImageNet dataset, using 20 inference steps. (a) FID of two models trained from scratch (b) FID of two models fine-tuned on Stable Diffusion v1.4.
@@ 4.3 Discussion
How much does image-noise distance reduce in the assignment? Tab. 4.3 shows the reduction in distance after the image-noise assignment. We find that the L2 distance only reduces by about 2%, with a slight increase observed at higher batch sizes. However, as shown in Fig. 3, the assignment is sufficient to effectively activate denoising at high noise levels, significantly boosting training efficiency, even though the distance change is low. We attribute the low distance reduction rate after the assignment to the extremely high dimensionality (3072 for each image of the CIFAR-10 dataset) of the image and noise space.
How much time does image-noise assignment cost? In Tab. 4.3, we indicate that our assignment method does not introduce significant extra overhead due to our utilization of quantized assignment in our practical implementation. Even for a large batch size per GPU of 1024, our algorithm only brings in an additional 22.8 ms, demonstrating the potential of utilization for future applications.
Immiscible and OT: who dominates the training efficiency enhancement? Our Immiscible Diffusion claims to enhance training efficiency by improving miscibility in noisy diffusion steps. However, the method we take towards Immiscible Diffusion, i.e. linear assignment between image and noise, also serve as a roughly approximate OT between image and noise, which might intuitively benefit the diffusion through straightening the diffusion paths [36, 50]. However, the previous section shows that the image-noise distance is only reduced by ~2%, motivating us to ask if OT is really the dominant factor?
To answer this question, we ablate these two factors: OT and immiscibility. We design a non-OT Immiscible Diffusion experiment which keeps the immiscible property while not involving the OT. This is achieved by assigning images to the flipped noise whose all dimensions are reversed, while using the original noise during diffusion. In such a way, the image-noise pair no longer follows OT, but still qualifies the Immiscible Diffusion - i.e. images are still assigned to a limited area. Interestingly, we observe that the non-OT Immiscible Diffusion can still accelerate and enhance the diffusion training, which is nearly comparable to the OT Immiscible Diffusion in final stages, as shown in in Fig. 4.3. Considering that the non-OT version introduce miscibility in middle diffusion layers, which we posit for its difference to OT version, we conclude that Immiscible Diffusion is dominant in enhancing the diffusion model’s performance, compared to the benefits from OT.
Table 2: Image-noise data-point L2 distance reduction after the assignment for minimizing it and the time cost for the assignment.
@@
Figure 7: Ablation of OT in Immiscible Diffusion. FIDs of OT and non-OT Immiscible Diffusion indicates that it is the Immiscible Diffusion rather than OT that dominate the performance enhancement.
5 Conclusion, Limitations, and Future Work
Inspired by the immiscibility phenomenon in physics, we introduce Immiscible Diffusion, a method to improve image-noise mapping to accelerate diffusion training. Specifically, Immiscible Diffusion is an assignment-then-diffusion strategy. One way of it is to minimize the image-noise pair distance within a mini-batch so that each image is diffused to nearby noise areas. This simple approach requires only one line of code and includes a quantized-assignment strategy to reduce computational overhead.
Experiments show our Immiscible Diffusion approach speeds up Consistency Model’s training by approximately 3x on the CIFAR-10 dataset, 1.3x on the CelebA dataset, and 1.2x on the tiny-ImageNet dataset, as well as in conditional generation and fine-tuning on Stable Diffusion. Thus, we show that Immiscible Diffusion can generalize to across datasets, baselines and tasks. Further analysis is provided to explain how this works.
Limitation. The assignment strategy is one straightforward way for Immiscible Diffusion, but not necessarily optimal. Due to the limited computational resources, our experiments are mainly conducted on small-scale datasets, so we lack the validation on larger-scale datasets such as LAION. In future work, we will improve the assignment strategy to cater to practical utilization of conditional generation such as accelerating the general text-to-image or text-to-video diffusion training.
Broader impact. With the increased use of diffusion models for image and video generation, the training of diffusion models is certain to become an increasing portion of data center workloads. Moreover, training time is a significant bottleneck in model development. Our proposed method significantly improves the efficiency of diffusion model training. We believe that our method has the potential to accelerate progress and reduce the cost of development in this field.
Appendix A Supplemental Materials
A.1 Additional Experiment Results
A.1.1 Quantitative Training Efficiency Improvements for Immiscible Consistency Model
A.1.2 Quantitative FID Improvements for Immiscible DDIM with Different Inference Steps.
@@ A.1.1 Quantitative Training Efficiency Improvements for Immiscible Consistency Model
$\Delta$FID -0.15 -0.17 -0.88 -1.49 -2.57
A.1.3 Ablation on the Distance Measurement Methods in Noise Assignment.
We use the L2 norm for our experiments. However, we note that the L2 norm may face more challenges in distance evaluation in high-dimensional spaces compared to the L1 norm. Therefore, we compare the performance of immiscible DDIMs using assignments based on the L1 and L2 norms. The results, as illustrated in Tab. 5, show that using the L2 norm provides better performance than the L1 norm.
Table 5: FID of using L1 or L2 norm for noise assignment in immiscible DDIM on CIFAR-10.
  Training Steps (k) 70.2 93.6 117.0 140.4 163.8
DDIM 6.30 5.56 4.86 4.34 4.12
Immiscible DDIM using L2 Norm 5.28 4.56 4.13 3.81 3.70
Immiscible DDIM using L1 Norm 5.34 4.66 4.16 3.87 3.82
A.1.4 Ablation on the Batch Size on Immiscible DDIM.
The effectiveness of image-noise assignment can intuitively rely on the batch size. Therefore, we perform a comparison to see the effectiveness of Immiscible Diffusion on DDIM across a selected range of batch sizes, whose result is shown in Fig. 8. We observe that while larger batch sizes consistently accelerate the training as expected, its training efficiency enhancement is not as large as that from Immiscible Diffusion. Immiscible Diffusions continously improve the training efficiency and the performance in the whole selected range of batch sizes.
Figure 8: Effectiveness of Immiscible DDIM in a selected range of batch sizes.
A.2 Qualitative Evaluations of Immiscible Diffusion
A.2.1 Generated images from immiscible and baseline Consistency Models trained on CIFAR-10 (Top) and CelebA (Down) with the same training steps.
Figure 9: Qualitative comparison for Immiscible and baseline Consistency Model. We show images generated with the two models trained for 100k steps respectively. Compared to baseline method, immiscible models capture more details and more features of objects.
A.2.2 Generated images from immiscible and baseline Consistency Models trained on CIFAR-10 Dataset for 100k steps without cherry-picking.
@@ A.2 Qualitative Evaluations of Immiscible Diffusion
A.2.3 Generated images from immiscible and baseline Consistency Models trained on CelebA Dataset for 100k steps without cherry-picking.
@@
A.2.4 Generated images from immiscible and baseline stable diffusion models trained unconditionally on 10% ImageNet for 70k steps.
Figure 14: Images generated by immiscible and baseline Stable Diffusion trained unconditionally on ImageNet for 70k steps. We see that the Immiscible Stable Diffusion presents more reasonable modal and catch more general features and details.
A.2.5 Generated images from immiscible and baseline stable diffusion models trained unconditionally on 10% ImageNet Dataset for 70k steps without cherry-picking.
Figure 15: Generated images from immiscible and baseline stable diffusion models trained unconditionally on 10% ImageNet Dataset for 70k steps without cherry-picking
A.2.6 Generated images from immiscible and baseline stable diffusion models trained conditionally on ImageNet Dataset for 20k steps.
Figure 16: Generated images from immiscible and baseline stable diffusion models trained conditionally on ImageNet Dataset for 20k steps.